OpenStudy (lifeisadangerousgame):

A cellular phone company currently has three towers, one in Tampa, one in Orlando, and one in Gainesville to serve the central Florida region. Orlando is 80 miles east of Tampa and Gainesville is 100 miles north of Tampa, what is the distance from Orlando to Gainesville?

OpenStudy (anonymous):

I think it would be 128. 06 miles

OpenStudy (lifeisadangerousgame):

How?

OpenStudy (anonymous):

Pythagorean theroam. Basically this creates a right triangle, so a^2+b^2=c^2. a is Tampa to Orlando, b is Gainesvilles to tampa and c is tampa to gainesvilles

OpenStudy (anonymous):

100^2=10000. 80^2=6400. Add them they are 16400 and then find the sq. root of that and you get 128.06

OpenStudy (lifeisadangerousgame):

ohh, okay that makes sense! Thanks ^_^

OpenStudy (anonymous):

No problem :)